Parth (parthkohli):

How to derive this?\[a + a^2 + a^3\cdots a^n = {a(a^n -1) \over a - 1} \]

OpenStudy (anonymous):

multiply by \(a\)

Parth (parthkohli):

Everyone replying. Am I the only one who doesn't know this? lol

OpenStudy (anonymous):

also it is wrong

Parth (parthkohli):

Mathematica is not wrong!

OpenStudy (anonymous):

denominator should be \(a-1\)

Parth (parthkohli):

Sorry, yes.

Parth (parthkohli):

I never make miskaes. lol

OpenStudy (ash2326):

Let S represent the sum \[S=a+a^2+a^3+...a^n\] multiply this by a \[aS= a^2+a^3+a^4+.....a^{n+1}\] subtract S by aS \[(1-a)S=a+0+0....-a^{n-1}\] we have \[(1-a)S=a-a^{n+1}\] Can you find S now?

OpenStudy (anonymous):

what @ash2326 said

Parth (parthkohli):

Gee, how did you people become such geniuses?

Parth (parthkohli):

And this is very similar to the geometric series sum proof.